Calculate the heat required to convert 2kg of ice at -12°c to steam at 100°c. L for ice= 336000j/kg, L for steam= 2260000. C for ice = 2100j/kg/k


1
Expert's answer
2021-06-07T08:30:52-0400

Heat required to heat the ice from -12oC to the melting (0oC):

"Q_1=c_{ice}m\Delta{T}=2100J/kg\cdot{K}\times2kg\times(0\degree{C}-(-12\degree{C}))=50400J"


Heat required to melt the ice at 0oC:

"Q_2=L_{ice}m=336000J/kg\times2kg=672000J"


Heat required to heat the water from 0oC to the boiling (100oC):

"Q_3=c_{water}m\Delta{T}=4184J/kg\cdot{K}\times2kg\times(100\degree{C}-0\degree{C})=836800J"


Heat required to evaporate the water at 100oC:

"Q_4=L_{steam}m=2260000J/kg\times2kg=4520000J"


The total heat is the sum of the heats required for each step:

"Q=Q_1+Q_2+Q_3+Q_4=50400J+672000J+836800J+4520000J=6080000J=6080kJ"


Answer: 6080 kJ